OpenStudy (anonymous):

Write the first five terms of the geometric sequence where t1 = -3, and r = -1.

sam (.sam.):

\[\large T_n=a_1r^{n-1}\]

OpenStudy (anonymous):

N=?

sam (.sam.):

Since \(T_1=-3\), Second term will be \[T_2=(-3)(-1)^{2-1}=3\] Third term will be \(T_3\)

sam (.sam.):

n is the number of term you are using

sam (.sam.):

For term 3 n will be 3

sam (.sam.):

\(a_1=First~term\)

OpenStudy (anonymous):

t3=9?

OpenStudy (anonymous):

-3,3,9,15,21

sam (.sam.):

\[T_3=(-3)(-1)^{3-1} =?\\ \\ T_4=(-3)(-1)^{4-1}=?\]

OpenStudy (anonymous):

T3=9 T4=15

sam (.sam.):

Why would \(T_3=9?\)

OpenStudy (anonymous):

(-3)(-1)^3-=2

OpenStudy (anonymous):

Cause I thought if t1=-3 t2=3 t3 would=9 since the common difference is 6?

sam (.sam.):